From: NASA HQ/SpaceRef/NASA Watch
Posted: Tuesday, February 14, 2006
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2006) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.††
The crew performed the second part of the three-part semi-annual maintenance of the TVIS treadmill in the Service Module (SM) floor "pit".†† [After yesterday's work was completed well ahead of timeline (for which the crew received kudos), today's task focused on completion of the inspections (lubricating components and checking for loose fasteners), replacing a damaged wire rope of the clamp rope assembly, and returning TVIS to nominal configuration.† Activities included returning the belt to its original tension level, reinstalling structural components, verifying torque on accessible fasteners, replacing the TVIS in its SM pit, and performing an unmanned 10-minute speed characterization test, which the CDR completed ahead of timeline. Ground personnel are presently analyzing the resulting data.† Exercise test data remain to be downlinked. Ground personnel will advise on return to normal operation when their analyses are complete.† A few anomalies were noticed during the maintenance activities, but none are expected to impact normal operation.† Assuming favorable engineering analysis, both crewmembers are to resume the bungee/eyebolt harness configuration used prior to this TVIS maintenance.]
FE Tokarev conducted the first stress test & saliva sampling of ESA's "Immuno" experiment, using its Saliva-Immuno Kit.† Valery's second stress test plus saliva sampling and data entry in the questionnaire will follow tonight before sleep time, after which the hardware will be stowed.†† [Objective of the experiment is to investigate immune neuro-endocrine reactions in the space environment by studying samples of saliva, blood and urine using collection kits and the biomedical (MBI) protection kit along with a stress-test questionnaire filled out by the subject.]
After yesterday's audit of water transfer equipment, the FE today worked on the temporarily installed KAV humidity condensate container #6577 in the SM, first drilling a 7-mm dia hole in the plastic cap at its bottom with the "Makita" power drill, then using this access point to transfer condensate samples (stemming from upstream of the FGS Gas-Liquid Mixture Filter of the SRVK-2M condensate water processor) to two empty drink bags for return to Earth.† Afterwards, he also collected two water samples from the EDV container of the SM's water supply system (SVO-ZV) in empty drink bags.
CDR McArthur ran another periodic atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen Sensor) and CDMK (CO2 Monitoring Kit).
Tokarev performed the regular weekly photo imagery of the BIO-5 Rasteniya-2 ("Plants-2") Lada-8 payload using the Nikon D1X digital camera with flash and copied all photos from the memory card to the RSK1 laptop for subsequent downlink to TsUP-Moscow via the BSR-TM telemetry channel.† [Rasteniya researches growth and development of plants under spaceflight conditions in the Lada-8 greenhouse.† The regular maintenance of the experiment (each Monday, Wednesday, Friday and Sunday) involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, topping off the water tank if ~20-25% of the total amount (4 liters) remains, and photo/video recording.]
The Flight Engineer completed the daily routine maintenance of the SM's environment control & life support system (SOZh), including its toilet system (ASU) and the daily updating/editing of the standard IMS "delta file", including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
Valery also conducted his regular checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at 20 degC, including a temperature check on its ART (automatic temperature recorder).†† [This daily monitoring/temp checking, carried on the Russian voluntary "time available" task list, will continue until 4/30.]
Both crewmembers completed their regular 2.5-hr. physical exercise program on the CEVIS cycle ergometer, RED resistive exerciser and VELO bike with bungee cord load trainer.†† [Valery's daily protocol for today required 1.5 hr on the RED, instead of the TVIS, and one hour on VELO plus load trainer (today: Day 4 of the first set).]
Afterwards, McArthur transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
At ~1:25pm EST, Bill held his weekly PFC (private family conference) via S-band/audio and Ku-band/MS-NetMeeting video.
New "Saturday Science" options for 2/18 were uplinked for the Science Officer for his selection later tonight.†† [The options are a review of FDI (Fluid Dynamics Investigation) activity of Saturday's CBOSS FDI (Cellular Biotechnology Support Systems-FDI) operations, or CBOSS Bubble Removal/Injection Rate Determination at the MWA (Maintenance Work Area). For 2/25, an EPO (Educational Payload Operation) demo of onboard food and sleep is being considered.]
On 2/11, ground personnel noticed issues with the KFX (Ku-band file transfer) software application on SSC-10 (Station Support Computer #10), which they could not access.† Yesterday, the crew swapped SSC-10's shell #1006 with shell #1004, but KFX remained inaccessible the ground today. Analysis is continuing.†† [The KFX application is used to transfer files from the ground to the onboard network.]
Update on SuitSAT (RadioSkaf):† The SuitSAT satellite continues to broadcast but the signal is reported to be extremely weak.† Hundreds of reports from individuals receiving the signal from all over the world have been logged and there has been tremendous interest in this project with nearly 4.5 million hits at the SuitSAT website (www.suitsat.org
Preparations are underway for an overnight "Campout" by the crew in the U.S. Airlock on 2/23-2/24 at 10.2 psia as an SDTO (Station Detailed Test Objective) test of this contingency.†† [Preps include the development of detailed Go/No-Go and termination criteria for the campout (which requires mask prebreathe), specifying unique conditions that would mandate termination of the test and re-opening of the Node starboard hatch.]
Today's CEO (Crew Earth Observation) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Internal waves, N Patagonian Shelf (excellent glint-enhanced views were expected just left of track on either side of the Valdes Peninsula), and Patagonian Glaciers (this pass had the best illumination and probably the best weather today. Trying for oblique, context views looking left of track towards the small Northern Patagonian Ice Field).
To date, more than 186,000 of CEO images have been taken in the first five years of the ISS, almost one third of the total number of images taken from orbit by astronauts.
† CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 12 crew visit:
Expedition 12 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Orbit† (as of this morning, 6:51am EST [= epoch]):
Events Ahead (all dates Eastern & tentative):
ISS Altitude History
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //